OpenStudy (anonymous):

Let x(t)=t-sin(t) and y(t)=1-cos(t) For t=5pi/6, find x(t) y(t) dx/dt dy/dt dy/dx tang lin slop

zepdrix (zepdrix):

Alright, so.... any trouble with the things listed before dy/dx? The dy/dx is going to be the most complicated of these. How bout the others?

OpenStudy (dumbcow):

\[\large \frac{dy}{dx} = \frac{\frac{dy}{dt}}{\frac{dx}{dt}}\]